What is the equation of a line that passes through the point (14,-9)and has a slope of Zero?
Answer by mangopeeler07(462)

y=mx+b

Plug in the slope for m
y=0x+b

Now plug in coordinates for y and x (14,-9)
-9=0(14)+b
-9=0+b
-9=b

Now plug in the values for m and b, but not x and y:
y=0x+-9

Now simplify
y=-9